Attention: Due to a bug in Google Chrome, this page may not function correctly. Click here to learn more 3. A one-tailed hypothesis test for a repeated-measures design Aa Aa E A cognitive psychologist is interested in whether increasing reading speed improves comprehension. He has students complete a reading comprehension task before and after a six-week speed-reading course In the beginning of the study, a randomly selected group of 81 students scored an average of 102 points on the reading comprehension task. Since the sample size is larger than 30, the cognitive psychologist can assume that the sampling distribution of Mo is normal. He plans to use a repeated-measures t-test. The cognitive psychologist identifies the null and alternative hypotheses as: Ho: HD Hi: HD Use the Distributions tool to find the critical region (s) for α-.05. The critical t-score, which is the value for t-scores that separates the tail(s) from the main body of the distribution and forms the critical region(s), is consider whether this is a one-tailed or two-tailed test.) (Hint: Remember to set the degrees of freedom on the tool and to Degrees of Freedom-95 esc F2 FS F7 4 IAI
Answer #1

Here we have to test whether increasing reading speed improves comprehension .i.e \muD > 0

So null and alternative hypothesis are

H​​​​​​0 : \muD\leq 0

H​​​​​​1​​​ : \muD > 0

t critical value at \alpha =0.05 and degrees of freedom =

d.f = n -1 = 81 -1 = 80

Here we have alternative hypothesis is right tailed

Using t table

t​​critical  = t​​​​​​0.95, 80 = 1.664

t​​​​critical = 1.664